### v1

@@ 4.1 Skills Development
Reflecting on their time on the DataSquad (or its immediate precursor [^9] Before the DataSquad was fully formalized, students were hired for the purposes of the DataSquad but not officially designated as members.), alumni reported substantial skill development across all eight assessed domains (Figure 1). Technical Documentation showed the highest average development (M=2.42), with 56% of alumni reporting maximum improvement. Data Wrangling/Cleaning followed closely (M=2.41), and Teamwork ranked third (M=2.28).
@@ 4.2 Career Influence
Overall, communication skills dominated career influence rankings, with three of the top four skills being interpersonal or communication-oriented rather than purely technical. This pattern suggests that while experience on the DataSquad facilitates the development of diverse capabilities, communication skills prove differentiating in professional contexts. Analysis of alumni rankings confirmed this trend: when asked "Drawing from your post-college experience, please choose the top 4 skill areas we should focus on for DataSquad member development." Technical Documentation appeared in 59.5% of responses, followed by Teamwork (46.0%) and Project Management (37.8%) (Figure 1). The similarity of these three charts indicates that the alumni felt that through their time on the DataSquad they gained skills they currently value in their careers and want current students to continue to get this experience.
@@ A.1 Learning Environment Ratings
The DataSquad learning environment received exceptionally positive ratings across all assessed dimensions (Figure 5). All five dimensions exceeded 88% “Yes” responses. Although the high ratings may reflect some level of response bias, the 100% encouragement rate is a positive sign, given that alumni who reported less positive feelings in the other categories agreed on the encouraging environment. It also suggests that there is some work to be done to improve the DataSquad experience for students who did not respond ’yes’ to some of the questions.
@@ A.2 Role Distribution and Multiple Pathways
Many students experienced multiple roles during their tenure (Figure 6). Among survey respondents, 43% served as Programmers/Data Scientists, 32% as Data Analysts, 22% as Technical Writers, and 11% as Project Management Interns. Importantly, 54% of alumni held two or more roles during their time on DataSquad, gaining breadth across the program’s offerings.
Figure 6: Distribution of roles held by alumni
@@ Appendix B Definitions
Confidence & Empowerment Feelings of accomplishment upon completion or participation in DataSquad projects
@@
Practical Data Experiences Real life experiences with messy data
@@
Across Experience Levels Communicating with clients and squad members of different tech backgrounds
Project Management Supervising a project from start to finish

### v2

@@ 4.1 Skills Development
Reflecting on their time on the DataSquad (or its immediate precursor [^9] Before the DataSquad was fully formalized, students were hired for the purposes of the DataSquad but not officially designated as members.), alumni reported substantial skill development across all eight assessed domains (Figure 1). Technical Documentation showed the highest average development (M=2.42), with 58% of alumni reporting maximum improvement. Data Wrangling/Cleaning followed closely (M=2.41), and Teamwork ranked third (M=2.28).
@@ 4.2 Career Influence
Overall, communication skills dominated career influence rankings, with three of the top four skills being interpersonal or communication-oriented rather than purely technical. This pattern suggests that while experience on the DataSquad facilitates the development of diverse capabilities, communication skills prove differentiating in professional contexts. Analysis of alumni rankings confirmed this trend: when asked "Drawing from your post-college experience, please choose the top 4 skill areas we should focus on for DataSquad member development." Technical Documentation appeared in 76% of responses, followed by Software Engineering (60%) and Teamwork (56%) (Figure 1). The similarity of these three charts indicates that the alumni felt that through their time on the DataSquad they gained skills they currently value in their careers and want current students to continue to get this experience.
@@ A.1 Learning Environment Ratings
The DataSquad learning environment received exceptionally positive ratings across all assessed dimensions (Figure 5). All five dimensions met or exceeded 88% “Yes” responses. Although the high ratings may reflect some level of response bias, the 100% encouragement rate is a positive sign, given that alumni who reported less positive feelings in the other categories agreed on the encouraging environment. It also suggests that there is some work to be done to improve the DataSquad experience for students who responded less favorably.
@@ A.2 Role Distribution and Multiple Pathways
Many students experienced multiple roles during their tenure (Figure 6). Among survey respondents, 57% served as Programmers/Data Scientists, 40% as Data Analysts, 27% as Technical Writers, and 13% as Project Management Interns. Importantly, approximately half of alumni held two or more roles during their time on DataSquad, gaining breadth across the program’s offerings.
Figure 6: Distribution of roles held by alumni. If multiple roles were held, both are counted and distinguished.
@@ Appendix B Definitions
Empowerment Feelings of accomplishment upon completion or participation in DataSquad projects
@@
Practical Data Real life experiences with messy data
@@
Experience Levels Communicating with clients and squad members of different tech backgrounds
Project Management Supervising and coordinating a project’s deliverables from start to finish
Documentation Collecting and recording data, documents, and other project information throughout the lifecycle of a project